OpenStudy (anonymous):

HI i need help i got a quad equation. that is like this Y=x^2 and under is 4 + 3x and uner is 4 + 5 and under is 2. now i'm suppose to be able to get find the y and x intercepts but am baffled by the facorisation am i suppose to take em out or work with them help :s

OpenStudy (amistre64):

can you draw out the problem to make it easier to parse?

OpenStudy (amistre64):

it sounds like youve got a fraction dividing a fraction ...

OpenStudy (amistre64):

\[\frac{a/b}{m/n}=\frac{a}{b}*\frac{n}{m}\]

OpenStudy (anonymous):